# Barvinok’s interpolation method meets Weitz’s correlation decay approach

### v1

@@ 1.1. The hard-core model
Barvinok’s interpolation method for the partition function of the hardcore model/the independence polynomial roughly works as follows. To approximate the polynomial $Z(G;\lambda)$ at some $0<\lambda<\lambda_{c}(\Delta)$, on the family, $\mathcal{G}_{\Delta}$, of graphs of maximum degree at most $\Delta$, one needs an open set $U\subset\mathbb{C}$ containing 00 and $\lambda$ such that for all $x\in U$ and $G\in\mathcal{G}_{\Delta}$ $Z(G;x)\neq 0.$ The existence of such a set is guaranteed by [PRSokal]. One then defines $f(x)=\log(Z(G;x))$ on $U$ (where we fix the principal branch of the logarithm) and uses the first logarithmically many (in the number of vertices) coefficients of the Taylor series of $f(x)$ to obtain a good approximation to $Z_{G}(\lambda)$. In case $U$ is a disk centered at zero of radius larger than $\lambda$, this follows from [Barbook]*Lemma 2.2.1, while if $U$ is of a different shape, for example a rectangle containing a real interval, this requires some additional work, see for example [Barbook]*Section 2.2.3 or [PatReg17]*Section 4.3. Computing the coefficients of $\log(Z_{G}(x))$ in a brute force manner leads to quasi-polynomial time algorithms. This has been improved to genuine polynomial time with a more refined algorithm in [PatReg17].
@@
At first sight these two algorithmic approaches may appear very different for both ingredients. However, recent developments have shown that the second ingredient for the respective methods (strong spatial mixing, resp. zero-freeness) are in fact closely related [LSSfisher, LSS2Delta, gamarnik2022correlation, regtsabsence].
@@
and it thus suffices to compute the first $\ell$ coefficients of the series expansion of $R_{G,v}(x)$ for each $v\in V$. We do this iteratively by expanding the ratio $R_{G,v}(x)$ in terms of ratios of smaller (induced subgraphs) just as Weitz does in his algorithm [Weitz]. In our opinion this provides a clear bridge between the algorithmic parts of the two approaches.
@@ 1.2. Our contributions
[…]
@@
We have to leave it as an interesting open question to devise an efficient algorithm for graphs of minimum degree at least $2$. We note that zeros of the independence polynomial of cycles approach $-1/4$ [ScottSokal, deboerratios], or in other words zeros of $Z_{\textrm{sfo}}(C_{n};t))$ approach $1/2$ (here $C_{n}$ denotes the cycle of length $n$). This implies that the interpolation method cannot be applied in a naive manner in case the graph has vertices of degree two. We note that [anand2025sinkfreeorientationslocalsampler] cannot handle graphs with degree two vertices for similar reasons (because the presence of zeros essentially implies that there is no constant lower bound on the marginal probability that a given vertex is not a sink).
@@
here $J$ denotes the $q\times q$ all ones matrix. By definition, we see that $H(G;1)=q^{-|V|}\hom(G,A)$. To obtain an efficient algorithm for approximating $H(G;A)$ for the family of graphs of maximum degree at most $\Delta$ based on the interpolation method one requires an open set $\mathcal{U}\subset\mathbb{C}$ that contains the points $x=0$ and $x=1$ such that $H(G;x)\neq 0$ for all $x\in\mathcal{U}$ and graphs $G$ of maximum degree at most $\Delta$. See [Barbook]*Section 2.2.3. For example, this is provided if $A\in\mathbb{C}^{q\times q}$ is close enough to the matrix $J$ in the infinity norm [Barbook] and also if $A=A(K_{q})$ provided $q\geq 1.998\Delta$ [bencs2024deterministicapproximatecountingcolorings, LSScorrelation]. Having such a zero-free region one then needs to compute the first $O(\log|V|/\varepsilon)$ coefficients of $\log(H(G,x))$. In [PatReg17] this is done expressing these coefficients as induced graph counts.
@@
Let $q\in\mathbb{N}_{\geq 2}$ and let $A$ be symmetric $q\times q$ matrix. Then there exists an algorithm that on input of a graph $G=(V,E)$ of maximum degree at most $\Delta$ and $m\in\mathbb{N}$ computes $\left(\log(H(G;x))\right)^{[m]}$ in time $O(e(q+1)\Delta)^{m}m^{3})$.
@@ 2. Preliminaries
Let $U\subseteq\mathbb{C}$ be an open set containing 00. For an analytic function $f:U\to\mathbb{C}$ with Taylor-series at 00, i.e.
@@
Let $f:U\to\mathbb{C}$ be an analytic function with a Taylor-series at 00 such that $f(0)=0$. Then there is an $O(m^{3})$-time algorithm that on input of $f^{[m]}(z)$ computes
@@
By definition we know that there is an analytic function $g(z)$ such $f(z)=zg(z)$. The Taylor series at $z=0$ of $1/(1-z)$ is given by $\sum_{k\geq 0}z^{k}$. Thus
@@ 3. The hard-core model
and let $R_{S,v}^{[k]}(x)$ be the $k$th Taylor polynomial of $R_{S,v}(x)$ around 00. Using this notation, we see that by (1.5),
@@
11:         $R_{i}\leftarrow R(G,S_{0},u_{i},k-1)$triangleright This is the recursive call
@@
19:     $p(x)\leftarrow x\left(\frac{1}{1+F(x)}\right)^{[k-1]}$triangleright Apply Lemma 2.1
@@
10:         $R_{v}(x)\leftarrow R(G,V(G),v,m)$triangleright Recursive call
11:         $F(x)=\left(\frac{1}{1+R_{v}(x)}\right)^{[m-1]}$triangleright Apply Lemma 2.1
@@ 4. Sink-free orientations
It is important to stress here that even though $U$ may not be equal to $V$ we do take the weight of vertex $v$ to be equal to $-t^{\deg_{G}(v)}$. As in the previous section it will be useful to consider ratios. To this end we define for a vertex $v\in U$ the following rational function,
@@ 4.1. A zero-free region
The next lemma establishes a zero-free disk for $Z_{\textrm{sfo},U}(G,t)$ allowing us to use Barinok’s interpolation method. In what follows for $r>0$, $B_{r}(0)$ denotes the disk of radius $r$ centered at 00. Let for a positive integer $\delta$,
@@ 4.2. The algorithm
So in case $\deg_{G}(v)>k$, we output 00. Note that it only takes $O(k)$ time to decide whether $\deg_{G}(v)>k$ and to list the set $\{u_{1},\ldots,u_{\ell}\}$ in case $\deg_{G}(v)\leq k$. In that case, to compute $R^{[k]}_{U,v}(t)$ we need to compute $F(t)$ and use Lemma 2.1 to compute $R^{[k]}_{U,v}$ (which takes time $O((k-\deg_{G}(v))^{3})$.) To compute $F(t)$ we need to call the algorithm to compute $R^{[k-\deg_{G}(v)]}_{U\setminus\{v,u_{1},\dots,u_{i-1}\},u_{i}}$ for $i=1,\ldots,\ell$ and multiply the resulting polynomials.
@@ 5. Chromatic Polynomial
Fix a total order $<$ of the edges of $G$. Recall that a set $F\subseteq E$ is called broken circuit free (abbreviated BCF) if $(V,F)$ is a forest (i.e. contains no cycles) and for each edge $e\notin F$ such that $F\cup e$ contains a cycle $C$, the edge $e$ is not the largest edge of $C$. Let $\mathcal{F}^{<}_{G}$ be the collection of BCF subgraphs of $G$. Then we know by Whitney’s theorem [Whitney] that
@@
Note that while $\mathcal{F}_{G}^{<}$ actually depends on the ordering $<$, but by definition, the polynomial $P(G;z)$ does not.
@@ 5.1. Enumeration of trees
Let $G=(V,E)$ be a graph of maximum degree $\Delta$ and $v$ to be a vertex of $G$. We assume that $E$ is equipped with a total order $<$. Let $\mathcal{Q}_{v}$ be the graph on the collection of subtrees of $G$ containing $v$, where two trees are connected if and only if they differ by exactly $1$ edge. Observe that in this graph the $m$th neighborhood of the single vertex tree $\{v\}$ is exactly the set of trees of at most $m$ edges. To do an efficient enumeration of this neighborhood in $\mathcal{Q}_{v}$, we will construct a spanning tree $\mathcal{T}_{v}$ of $\mathcal{Q}_{v}$ on which we can perform efficiently a graph search algorithm.
First let us describe the construction of $\mathcal{T}_{v}$. For each subtree $T\in V(\mathcal{Q}_{v})$ let us define a unique identifier: if $T$ has exactly $1$ edge $e$, then $ID(T)=(e)$. If $T$ consist of $m\geq 2$ edges, then $ID(T)=(e_{1},\dots,e_{m})$, where $e_{m}$ is the largest leaf edge of $T$, such that $v$ is not an isolated vertex of $T-e_{m}$ and $(e_{1},\dots,e_{m-1})=ID(T-e_{m})$. By definition $ID(T-e_{m})$ is exactly the sequence formed by the first $m-1$ elements of the sequence $ID(T)$; we call $T-e_{m}$ to be the parent of $T$. In particular, if $T^{\prime}$ is a parent of $T$, then $(T,T^{\prime})\in E(\mathcal{Q}_{v})$. Thus, if we only take the edges, that connects every tree with their parent, then we obtain a spanning tree $\mathcal{T}_{v}$ of $\mathcal{Q}_{v}$.
@@
  • computes $\{ID(T^{\prime})\leavevmode\nobreak\ |\leavevmode\nobreak\ T^{\prime}\in\mathcal{Q}_{v},\textrm{$T$ is the parent of $T^{\prime}$}\}$ in time $O(\Delta|T|)$.
@@
For a graph $G=(V,E)$ of with a given total ordering of its edges $<$ and a forest $F\subset E$ we call an edge $f\in E\setminus F$ broken if $F\cup f$ contains a cycle $C$ and $f$ is the largest edge in that cycle. We denote the collection of broken edges of $F$ by $B(F).$
@@ 6. Graph homomorphisms
We first need a few definitions to realize $H(G,x)$ as a forest generating function. Consider a total ordering $<$ on the edges of $G$. Let for a tree $T\subseteq E$

### v2

@@ 1.1. The hard-core model
Barvinok’s interpolation method for the partition function of the hardcore model/the independence polynomial roughly works as follows. To approximate the polynomial $Z(G;\lambda)$ at some $0<\lambda<\lambda_{c}(\Delta)$, on the family, $\mathcal{G}_{\Delta}$, of graphs of maximum degree at most $\Delta$, one needs an open set $U\subset\mathbb{C}$ containing 0 and $\lambda$ such that for all $x\in U$ and $G\in\mathcal{G}_{\Delta}$ $Z(G;x)\neq 0.$ The existence of such a set is guaranteed by [PRSokal]. One then defines $f(x)=\log(Z(G;x))$ on $U$ (where we fix the principal branch of the logarithm) and uses the first logarithmically many (in the number of vertices) coefficients of the Taylor series of $f(x)$ to obtain a good approximation to $Z_{G}(\lambda)$. In case $U$ is a disk centered at zero of radius larger than $\lambda$, this follows from [Barbook]*Lemma 2.2.1, while if $U$ is of a different shape, for example a rectangle containing a real interval, this requires some additional work, see for example [Barbook]*Section 2.2.3 or [PatReg17]*Section 4.3. Computing the coefficients of $\log(Z_{G}(x))$ in a brute force manner leads to quasi-polynomial time algorithms. This has been improved to genuine polynomial time with a more refined algorithm in [PatReg17].
@@
At first sight these two algorithmic approaches may appear very different for both ingredients. However, recent developments have shown that the second ingredient for the respective methods (strong spatial mixing, resp. zero-freeness) are in fact closely related [LSSfisher, LSS2Delta, Shaounified, gamarnik2022correlation, regtsabsence, shao2024zero].
@@
and it thus suffices to compute the first $\ell$ coefficients of the series expansion of $R_{G,v}(x)$ for each $v\in V$. We do this iteratively by expanding the ratio $R_{G,v}(x)$ in terms of ratios of smaller (induced subgraphs) just as Weitz does in his algorithm [Weitz]. In our opinion this provides a clear bridge between the algorithmic parts of the two approaches [^2] We note that some time after the initial posting of the present paper to the arXiv, Shao and Shi [shao2025zero] showed a way of using zero-freeness in combination with essentially the first ingredient of Weitz’s approach to design efficient approximation algorithms for evaluating partition functions, that relies on similar ideas as those that have been developed here..
@@ 1.2. Our contributions
and realize that $Z_{\textrm{sfo}}(G;1/2)=2^{-|E|}\textrm{sfo(G)}$. We prove in Section 4 a suitable zero-free region for this polynomial, providing the first ingredient for Barvinok’s interpolation method. For the second ingredient we extend our algorithm for computing Taylor coefficients of $\log(Z(G;x))$ to $\log(Z_{\textrm{sfo}}(G;t))$ and thereby we obtain the following result:
@@
We have to leave it as an interesting open question to devise an efficient algorithm for graphs of minimum degree at least $2$. We note that zeros of the independence polynomial of cycles approach 1/4-1/4 [ScottSokal, deboerratios], or in other words zeros of $Z_{\textrm{sfo}}(C_{n};t))$ approach $1/2$ (here $C_{n}$ denotes the cycle of length $n$). This implies that the interpolation method cannot be applied in a naive manner in case the graph has vertices of degree two. We note that [anand2025sinkfreeorientationslocalsampler] cannot handle graphs with degree two vertices for similar reasons (because the presence of zeros essentially implies that there is no constant lower bound on the marginal probability that a given vertex is not a sink).
@@
here $J$ denotes the $q\times q$ all ones matrix. By definition, we see that $H(G;1)=q^{-|V|}\hom(G,A)$. To obtain an efficient algorithm for approximating $H(G;A)$ for the family of graphs of maximum degree at most $\Delta$ based on the interpolation method one requires an open set $\mathcal{U}\subset\mathbb{C}$ that contains the points $x=0$ and $x=1$ such that $H(G;x)\neq 0$ for all $x\in\mathcal{U}$ and graphs $G$ of maximum degree at most $\Delta$. See [Barbook]*Section 2.2 and [PatReg17]. For example, this is provided if $A\in\mathbb{C}^{q\times q}$ is close enough to the matrix $J$ in the infinity norm [Barbook] and also if $A=A(K_{q})$ provided $q\geq 1.998\Delta$ [bencs2024deterministicapproximatecountingcolorings, LSScorrelation]. Having such a zero-free region one then needs to compute the first $O(\log|V|/\varepsilon)$ coefficients of $\log(H(G,x))$. In [PatReg17] this is done expressing these coefficients as induced graph counts.
@@
Let $q\in\mathbb{N}_{\geq 2}$ and let $A$ be symmetric $q\times q$ matrix. Then there exists an algorithm that on input of a graph $G=(V,E)$ of maximum degree at most $\Delta$ and $m\in\mathbb{N}$ computes $\left(\log(H(G;x))\right)^{[m]}$ in time $O((e(q+1)\Delta)^{m}m^{3})$.
@@ 2. Preliminaries
Let $U\subseteq\mathbb{C}$ be an open set containing 0. For an analytic function $f:U\to\mathbb{C}$ with Taylor-series at 0, i.e.
@@
Let $f:U\to\mathbb{C}$ be an analytic function with a Taylor-series at 0 such that $f(0)=0$. Then there is an $O(m^{3})$-time algorithm that on input of $f^{[m]}(z)$ computes
@@
By definition we know that there is an analytic function $g(z)$ such $f(z)=zg(z)$. The Taylor series at $z=0$ of 1/(1z)1/(1-z) is given by $\sum_{k\geq 0}z^{k}$. Thus
@@ 3. The hard-core model
and let $R_{S,v}^{[k]}(x)$ be the $k$th Taylor polynomial of $R_{S,v}(x)$ around 0. Using this notation, we see that by (1.5),
@@
11:   $R_{i}\leftarrow R(G,S_{0},u_{i},k-1)$ $\triangleright$ This is the recursive call
@@
19:  $p(x)\leftarrow x\left(\frac{1}{1+F(x)}\right)^{[k-1]}$ $\triangleright$ Apply Lemma 2.1
@@
10:   $R_{v}(x)\leftarrow R(G,V(G),v,m)$ $\triangleright$ Recursive call
11:   $F(x)=\left(\frac{1}{1+R_{v}(x)}\right)^{[m-1]}$ $\triangleright$ Apply Lemma 2.1
@@ 4. Sink-free orientations
It is important to stress here that even though $U$ may not be equal to $V$ we do take the weight of vertex $v$ to be equal to $t^{\deg_{G}(v)}$. As in the previous section it will be useful to consider ratios. To this end we define for a vertex $v\in U$ the following rational function,
@@ 4.1. A zero-free region
The next lemma establishes a zero-free disk for $Z_{\textrm{sfo},U}(G,t)$ allowing us to use Barvinok’s interpolation method. In what follows for $r>0$, $B_{r}(0)$ denotes the disk of radius $r$ centered at 0. Let for a positive integer $\delta$,
@@ 4.2. The algorithm
So in case $\deg_{G}(v)>k$, we output 0. Note that it only takes $O(k)$ time to decide whether $\deg_{G}(v)>k$ and to list the set $\{u_{1},\ldots,u_{\ell}\}$ in case $\deg_{G}(v)\leq k$. In that case, to compute $R^{[k]}_{U,v}(t)$ we need to compute $F(t)$ and use Lemma 2.1 to compute $R^{[k]}_{U,v}$ (which takes time $O((k-\deg_{G}(v))^{3})$.) To compute $F(t)$ we need to call the algorithm to compute $R^{[k-\deg_{G}(v)]}_{U\setminus\{v,u_{1},\dots,u_{i-1}\},u_{i}}$ for $i=1,\ldots,\ell$ and multiply the resulting polynomials.
@@ 5. Chromatic Polynomial
Fix a total order << of the edges of $G$. Recall that a set $F\subseteq E$ is called broken circuit free (abbreviated BCF) if $(V,F)$ is a forest (i.e. contains no cycles) and for each edge $e\notin F$ such that $F\cup e$ contains a cycle $C$, the edge $e$ is not the largest edge of $C$. Let $\mathcal{F}^{<}_{G}$ be the collection of BCF subgraphs of $G$. Then we know by Whitney’s theorem [Whitney] that
@@
Note that while $\mathcal{F}_{G}^{<}$ actually depends on the ordering <<, but by definition, the polynomial $P(G;z)$ does not.
@@ 5.1. Enumeration of trees
Let $G=(V,E)$ be a graph of maximum degree $\Delta$ and $v$ to be a vertex of $G$. We assume that $E$ is equipped with a total order <<. Let $\mathcal{Q}_{v}$ be the graph on the collection of subtrees of $G$ containing $v$, where two trees are connected if and only if they differ by exactly $1$ edge. Observe that in this graph the $m$th neighborhood of the single vertex tree $\{v\}$ is exactly the set of trees of at most $m$ edges. To do an efficient enumeration of this neighborhood in $\mathcal{Q}_{v}$, we will construct a spanning tree $\mathcal{T}_{v}$ of $\mathcal{Q}_{v}$ on which we can perform efficiently a graph search algorithm.
First let us describe the construction of $\mathcal{T}_{v}$. For each subtree $T\in V(\mathcal{Q}_{v})$ let us define a unique identifier: if $T$ has exactly $1$ edge $e$, then $ID(T)=(e)$. If $T$ consist of $m\geq 2$ edges, then $ID(T)=(e_{1},\dots,e_{m})$, where $e_{m}$ is the largest leaf edge of $T$, such that $v$ is not an isolated vertex of $T-e_{m}$ and $(e_{1},\dots,e_{m-1})=ID(T-e_{m})$. By definition $ID(T-e_{m})$ is exactly the sequence formed by the first m1m-1 elements of the sequence $ID(T)$; we call $T-e_{m}$ to be the parent of $T$. In particular, if $T^{\prime}$ is a parent of $T$, then $(T,T^{\prime})\in E(\mathcal{Q}_{v})$. Thus, if we only take the edges, that connects every tree with their parent, then we obtain a spanning tree $\mathcal{T}_{v}$ of $\mathcal{Q}_{v}$.
@@
  • computes $\{ID(T^{\prime})\penalty 10000\ |\penalty 10000\ T^{\prime}\in\mathcal{Q}_{v},\textrm{$T$ is the parent of $T^{\prime}$}\}$ in time $O(\Delta|T|)$.
@@
For a graph $G=(V,E)$ of with a given total ordering of its edges << and a forest $F\subset E$ we call an edge $f\in E\setminus F$ broken if $F\cup f$ contains a cycle $C$ and $f$ is the largest edge in that cycle. We denote the collection of broken edges of $F$ by $B(F).$
@@ 6. Graph homomorphisms
We first need a few definitions to realize $H(G,x)$ as a forest generating function. Consider a total ordering << on the edges of $G$. Let for a tree $T\subseteq E$